# IDDM: Bridging Synthetic-to-Real Domain Gap from Physics-Guided Diffusion for Real-world Image Dehazing

### v1

@@ 5.3 Ablation Study
Studies of our framework components: Due to the intricate interdependencies within our training process, it is not feasible to present ablation results in a simple additive form (e.g., A, A+B, A+B+C). Therefore, our ablation study is organized into several comparative groups: 1) w/o IDDM, which only use the normal diffusion model DDPM for image dehazing; 2) w/o $\mathbf{h}_{t}$, which means do not utilize $\mathbf{h}_{t}$ for noise predictor’s training and sampling; 3) w/o stage 3, which means use the denoising Unet checkpoint trained at stage 1; As shown in Fig. 6 (b), without IDDM, the image remains hazy and unclear. In Fig. 6 (c), the absence of $\mathbf{h_{t}}$ results in the color bias and distortion of the recovery image. Fig. 6 (d) shows that our model enhances contrast and color at Stage 3. In Fig. 6 (e) and (f), the results trained with both indoor and outdoor datasets are provided, which shows that IDDM is robust for multi-domain generalization problems. We also present qualitative results with all models trained on OTS, shown in Table 3.
Hyperparameter Study: On the BEDDE dataset, we conducted comprehensive ablation studies on key hyperparameters while maintaining constant values for other variables ($T=1000$, $\beta\in[0.4,1.5]$, and $A\in[0.7,1.0]$). As shown in Table 4 and Table 5, the experimental results demonstrate that the selected hyperparameters are optimal.
@@ 6 Conclusion and Discussion
In this paper, we present Image Dehazing Diffusion Models (IDDM), a novel approach addressing the challenge of synthetic-to-real domain discrepancy in image dehazing. By integrating the atmospheric scattering model into diffusion frameworks, our method learns the physical process of haze formation during the forward diffusion phase, enabling effective dehazing through the sampling process.
IDDM bridges the gap between synthetic training data and real-world applications by explicitly modeling how haze physically develops and interacts with scene elements. Extensive experiments demonstrate that our approach consistently outperforms state-of-the-art methods across three real-world haze datasets. Notably, even when trained solely on the ITS indoor synthetic dataset, which differs significantly from real-world outdoor scenes, our model achieves impressive dehazing results. This cross-domain robustness stems directly from our incorporation of physical principles, which prevents artifacts and enables effective generalization across vastly different visual domains. Our results confirm that by teaching diffusion models to understand the physics of haze formation, IDDM successfully overcomes the synthetic-to-real domain gap that has long challenged image dehazing research.

### v3

@@ 5.3 Ablation Study
Studies of our framework components: Due to the intricate interdependencies within our training process, it is not feasible to present ablation results in a simple additive form (e.g., A, A+B, A+B+C). Therefore, our ablation study is organized into several comparative groups: 1) w/o IDDM, which only uses the normal diffusion model DDPM for image dehazing; 2) w/o $\mathbf{h}_{t}$, which means that you do not utilize $\mathbf{h}_{t}$ for noise predictor’s training and sampling; 3) w/o Eq. 14, which means directly applying simple normalization after sampling. 4) The results use indoor synthetic dataset (ITS) or outdoor synthetic dataset (OTS) for training.
As shown in Fig. 6 (b), without IDDM, the image remains hazy and unclear. In Fig. 6 (c), the absence of $\mathbf{h}_{t}$ results in the color bias and distortion of the recovery image. In Fig. 6 (e) and (f), the results trained with both indoor and outdoor datasets are provided, which shows that IDDM is robust for multi-domain generalization problems. We observe that in 6 (e) and (f), the model trained on the ITS dataset removes more distant haze than that trained on the OTS dataset. This is because indoor environments often do not contain haze, whereas in natural scenes, a certain amount of haze is always present, and this cannot be avoided even in clear outdoor datasets. We also present qualitative results on RTTS [reside] with all models trained on OTS, shown in Table 4.
Hyperparameter Study: On the BEDDE dataset, we conducted comprehensive ablation studies on key hyperparameters while maintaining constant values for other variables ($T=1000$, $\sigma\in[0.4,1.5]$, and $A\in[0.7,1.0]$). As shown in Table 5 and Table 6, the experimental results demonstrate that the selected hyperparameters are optimal.
@@ 6 Conclusion and Discussion
In this paper, we present Image Dehazing Diffusion Models (IDDM), a novel approach addressing the challenge of synthetic-to-real domain discrepancy in image dehazing. By integrating the atmospheric scattering model into diffusion frameworks, our method learns the physical process of haze formation during forward diffusion, enabling effective dehazing through sampling.
Extensive experiments demonstrate that IDDM consistently outperforms state-of-the-art methods across three real-world haze datasets. Notably, even when trained solely on ITS indoor synthetic dataset, our model achieves impressive dehazing results on real-world outdoor scenes. This cross-domain robustness stems from our incorporation of physical principles, which prevents artifacts and enables effective generalization across different visual domains.